A bank loans a total of $7300 to two businesses. The bank charges the first business 5% annual interest, and charges the second business 8% annual interest. At the end of a year, the businesses pay the bank $440 in interest. How much did the bank loan each business?

Answer by checkley71(8403)

.05X+.08(7300-X)=440 OR .O5X+584-.08X=440 OR -.03X=-144 OR .03X=144 OR
X=144/.03 OR X=4800 THEN .05*4800=240 & 7300-4800=2500 OR 2500*.08=200
THEN 240+200=450
SO THE BUSINESSES RECEIVED LOANS OF $4800 @ 5% & $2500 @ 8%